OpenStudy (anonymous):

prove the statement true : 1^2 +4^2 + 7^2 + ..... + (3n-2)^2 = [n (6n^2 - 3n -1)] / 2

myininaya (myininaya):

did you see my earlier reply to this?

OpenStudy (anonymous):

i dont see it :(

myininaya (myininaya):

oh wait nvm ignore the early one do you know anything about induction?

OpenStudy (anonymous):

some1 tell me put n=1 and see both side equal , then that's proved

OpenStudy (anonymous):

but when i try another n=2 then it s not equal

myininaya (myininaya):

for n=1 we have (3*1-2)^2=1^2=1 on lhs we have 1*(6*1^2-3*1-1)/2=2/2=1 on rhs so now we get to assume is true for k>=1 and now we have to show it is true for k+1

myininaya (myininaya):

1^2+4^2+7^2+....+(3k-2)^2+(3(k+1)-2)^2 we want to show this equals the other side (whereever this is n we need k+1 to be there)

myininaya (myininaya):

so we know 1^2+4^2+...+(3k-2)^2=k(6k^2-3k-1)/2 so we have k[6k^2-3k-1]/2+(3(k+1)-2)^2 we need to manipulate this so that it is (k+1)*(6(k+1)^2-3(k+1)-1)/2 ok? try to do this

myininaya (myininaya):

oh hello, n=2 it does work for n=2 we have on lhs 1^2+4^2=1+16=17 on rhs we have 2(6*2^2-3*2-1)/2=24-6-1=17 so does work for n=2

OpenStudy (anonymous):

it wont work for the k+1 because the LHS ^ 2 and the RHS ^ 3 ? So can you explain more about this ? :-<

myininaya (myininaya):

?

myininaya (myininaya):

what do you mean?

OpenStudy (anonymous):

let me try this again

myininaya (myininaya):

here i will scan what i did. you really have to know some algbra for this

myininaya (myininaya):

i added in alot of zeroes i forced it to be what i want and then the left overs canceled with the other left overs :)

myininaya (myininaya):

do you have any questions?

OpenStudy (anonymous):

i got it now really appreciate for your help

myininaya (myininaya):

so you actually understood all the weird stuff i did? lol

OpenStudy (anonymous):

no i did it another way and got the answer

myininaya (myininaya):

if i looked back on this, i would be like wtf

OpenStudy (anonymous):

LOLLLLL

myininaya (myininaya):

lol i was trying to do it another way and i kept failing

myininaya (myininaya):

can you post what you did

OpenStudy (anonymous):

i dont know how to post the paper attaach in here , and i dont have the scan machine :(

OpenStudy (anonymous):

but i will try to type it then

myininaya (myininaya):

no worries lol